CAVITY RINGDOWN ABSORPTION SPECTRUM OF THE T_{1}(n, π^{∗}) ← S_{0} TRANSITION OF 2-CYCLOHEXEN-1-ONE

KATHERINE L. ZABRONSKY, MICHAEL O. McANALLY, DANIEL J. STUPCA, NATHAN R. PILLSBURY, and \underline{STEPHEN DRUCKER}, Department of Chemistry, University of Wisconsin-Eau Claire, Eau Claire, WI 54702.

The cavity ringdown (CRD) absorption spectrum of 2-cyclohexen-1-one (CHO) was recorded over the range 401.5–410.5 nm in a room-temperature gas cell. The very weak band system (ϵ ≤ 0.02 dm^{3} mol^{−1} cm^{−1}) in this region is due to the T_{1}(n, π*) ← S_{0} electronic transition. The 0_{0}^{0} origin band was assigned to the feature observed at 24,558.6 ±0.3 cm^{−1}. We have assigned about 25 vibronic transitions in a region extending from −200 to +350 cm^{−1} relative to the origin band. From these assignments we determined fundamental frequencies for several vibrational modes in the T_{1} excited state. The table below compares their frequencies to corresponding values measured for CHO vapor in the S_{0} electronic ground state (via far-IR spectroscopy)^{a} and the S_{1}(n, π*) excited state (via near-UV CRD spectroscopy).^{b}

Low-frequency fundamentals (cm^{−1}) of CHO vapor

Mode  Description  S_{0}  S_{1}(n, π^{∗}) T_{1}(n, π^{∗})
39  ring twist  99.2  122.1  99.5
38  bend (inversion of C-5)  247  251.9  253.2
37  C=C twist  304.1  303.3  247.8
36  C=O wag  485  343.9  345.5

For ν_{39} and ν_{37}, the differences between S_{1} and T_{1} frequencies are noteworthy. These differences suggest that the electron delocalization associated with the π^{∗} ← n chromophore in CHO is substantially different for singlet vs. triplet excitation.

aT. L. Smithson and H. Wieser, J. Chem. Phys. 73, 2518 (1980); M. Z. M. Rishard and J. Laane, J. Molec. Struct. 976, 56 (2010).

bM. Z. M. Rishard, E. A. Brown, L. K. Ausman, S. Drucker and J. Laane, J. Phys. Chem. A 112, 38 (2008).